lim x ➡0 (lim y ➡0 for ysinx/|x|) does not exist. Say true or false


1
Expert's answer
2022-01-24T17:04:28-0500
"\lim_{x \rightarrow 0 } \bigg(\lim_{y \rightarrow 0 } \frac{y \sin x}{|x|}\bigg)\\
\lim_{x \rightarrow 0 } \bigg(\lim_{y \rightarrow 0 } \frac{0 \cdot \sin x}{|x|}\bigg)\\
\lim_{x \rightarrow 0 } \bigg(\lim_{y \rightarrow 0 } \frac{0 }{|x|}\bigg)\\
\lim_{x \rightarrow 0 } \bigg(\lim_{y \rightarrow 0 } 0 \bigg)\\
\lim_{x \rightarrow 0 } \big(0 \big)\\
\lim_{x \rightarrow 0 } 0 =0\\"

Therefore,


"\lim_{x \rightarrow 0 } \bigg(\lim_{y \rightarrow 0 } \frac{y \sin x}{|x|}\bigg) =0\\"

and thus exist (FALSE)